OpenStudy (anonymous):

What is the equation of the line that has a slope of -3 and passes through (1,6) A.y=-3x-1 B.y=-3x+11 C.y=-3x+9 D.y=-3x+1

OpenStudy (anonymous):

You are given the slope so plug that into the formula: y-y1=-3(x-x1) to get your equation

OpenStudy (anonymous):

@Juanita35

OpenStudy (anonymous):

So i plug in the point i have

OpenStudy (anonymous):

yes, so you plug in the point and solve for the equation

OpenStudy (anonymous):

what did you get?

OpenStudy (anonymous):

So it would be y-6=-3(1-x1

OpenStudy (anonymous):

almost

OpenStudy (anonymous):

y - 6 = -3(x - 1)

OpenStudy (anonymous):

and then you would have y - 6 = -3x + 3

OpenStudy (anonymous):

Oh i see you distrubated

OpenStudy (anonymous):

you want to distribute

OpenStudy (anonymous):

and get y all by itself

OpenStudy (anonymous):

By moving the 6

OpenStudy (anonymous):

correct, so you would add 6 to both sides

OpenStudy (anonymous):

Would it be -3x+9

OpenStudy (anonymous):

nice job :)

OpenStudy (anonymous):

Thank you so much !

OpenStudy (anonymous):

I am glad I can help